ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 218, 2017

Arakelian, N. and H. Borges Frobenius nonclassicality of Fermat curves with respect to cubics

273

Astashkin, S. Rademacher functions in weighted symmetric spaces  371

Borges, H. See N. Arakelian Brandenberg, R. and B. Gonzalez Merino The asymmetry of

complete and constant width bodies in general normed spaces and the Jung constant

489

Chaudouard, P. Sur la contribution unipotente dans la formule des traces d’Arthur pour les groupes généraux linéaires

175

Cibulka, J., M. Korbelář, J. Kynčl, V. Mészáros, R. Stolař and P. Valtr On three measures of non-convexity

331

Conley, C. T. and B. D. Miller Orthogonal measures and ergodicity

83

Dèbes, P. On the Malle conjecture and the self-twisted cover  101

Doucha, M. Universal and homogeneous structures on the Urysohn and Gurarij spaces

299

Eisner, I. Exotic cluster structures on SL_{n} with Belavin–Drinfeld data of minimal size, I. The structure

391

Eisner, I. Exotic cluster structures on SL_{n} with Belavin–Drinfeld data of minimal size, II. Correspondence between cluster structures and Belavin–Drinfeld triples

445

Gonzalez Merino, B. See R. Brandenberg Hytönen, T. P., L. Roncal and O. Tapiola Quantitative

weighted estimates for rough homogeneous singular integrals

133

Kalaj, D. On J. C. C. Nitsche type inequality for annuli on Riemann surfaces

67

Komjáth, P. and S. Shelah Consistently P(ω_{1}) is the union of less than 2^{ℵ_{1}} strongly independent families

165

Korbelář,M. See J. Cibulka Kynčl, J. See J. Cibulka Mészáros, V. See J. Cibulka Miller, B. D. See C. T. Conley Roncal, L. See T. P. Hytönen Schmüdgen, K. A general fibre theorem for moment problems and

some applications 43 Shelah, S. See P. Komjáth

Simon, P. VC-sets and generic compact domination  27

Stolař, R. See J. Cibulka Tapiola, O. See T. P. Hytönen Valtr, P. See J. Cibulka Wang, J. Rank three Nichols algebras of diagonal type over arbitrary

fields 1